OpenStudy (dallaslifebaby01):

What is the solution to -2(3x - 9) + 5x = -10?

OpenStudy (haseeb96):

first multiply -2 with the parenthesis

OpenStudy (dallaslifebaby01):

the kids I'm sitting have done that already I'm truing to check their answer

OpenStudy (dallaslifebaby01):

@Haseeb96

OpenStudy (haseeb96):

okay what you got ?

OpenStudy (dallaslifebaby01):

-28 -8 8 28 these are the answer choices and she got -8

OpenStudy (haseeb96):

wrong let me solve it -6x + 18 + 5x =-10 -x = -10 -18 x = 28 28 is the correct answer

OpenStudy (dallaslifebaby01):

okay I'm going to show her where she went wrong

OpenStudy (haseeb96):

okay